1. Find a general solution of X = AX with -5 4 (1) A= (2) A= 4 -5 2 3 1 -2 4 (3) A= 2 2 3 -3 (4) A= -6 2 4
Answer #1

E A = -5 Now find eigenvalue and eigenvector of A - IA-211-0 1 1-5-d 4 16 = (-5-4) -1670 (4+5) = d+5 = 24 - 1 = -2, -1. - eig: eigenvector corresponding eigenvalue dz=-1 is -L : The general solution. x = a é 1941;] et[] + с. е Ae [132] - - 2 Now find4-6-1-(7 41 68. [ 17**** [ [1] [3]-[:] .24-22=0 2 = 2 [ma] [m] => [!] eigenvector corresponding eigenvalue diaries -- [1] ..diso -2 R₂-3 R₂ R2 276UR, A-(OI 325R [. :] 3 [B]=[-][- :) : mn2-0 22 =22 - [ ] [ ] = z2[:] eigenvector corresponding eigenvalA= 2 Now find eigenvalue and eigenvector of A. ; IA-d11=0 1-6-2 4 2 -4-2 >(-6-4) (-4-22 -820 2 +10d +24-820 22 +10d + 16-0 ->N N (1) eigenvector corresponding eigenualue dz=r2 is 01 the general solution <=6€*[*3] +60=4[!]